A Complete Bibliography of Publications in the ACM Symposia on Theory of Computing (STOC) for

1960–1969

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 April 2013 Version 1.01

Title word cross-reference

Abstract [BG69, Cha69, Coo69, Ros69, Sav69]. algebras [She69]. algorithm [Ull69]. algorithms [Blu69b]. ambiguity [Ull69]. analysis [PR69]. arbitrary [HI69]. arithmetic [Avi69]. arrays [ALY69]. automata [Tha69].

boolean [Spi69]. bounds [MM69].

California [ACM69]. certain [Spi69]. changing [You69]. Classes [MM69, Bas69, Spi69]. compilers [Blu69a]. complexity [Avi69, Hod69, Lov69]. Computability [HI69]. computable [Bas69, MM69]. computation [Col69, MM69, Spi69]. computational [Avi69]. Computing [ACM69]. Conference [ACM69]. constituent [PR69]. construction [Kor69]. Context [PR69, Rou69, AU69, Ull69]. Context-free [Rou69, PR69]. Context-sensitive [PR69].

defined [MP69, MM69]. del [ACM69]. Detailed [Coo69, Sav69]. Deterministic [Sav69, Cha69, RS69]. down [RS69]. effective [Blu69b]. Efficient [Kor69]. enumerated [You69]. expressions [Joh69]. Extended [BG69].

families [Cha69, Ros69]. features [Zei69]. fields [HI69]. finite [Tha69]. Formal [Zei69]. Formalization [MP69]. framework [ALY69]. Free [Joh69, AU69, PR69, Rou69, Ull69]. functions [Avi69, Bas69, MP69, MM69, Spi69].

general [She69]. generalized [Tha69]. geometric [Hod69]. grammar [AU69]. grammars [RS69, Rou69]. groups [Joh69].

immediate [PR69]. inherent [Ull69]. Intercalation [Ogd69]. iterative [ALY69].

Languages [She69, Blu69a, BG69, Cha69, Fis69, Ogd69, PR69, Ull69, Zei69]. logical [Hod69]. LR [Kor69].

machines [ALY69, Col69, Coo69, Sav69]. Marina [ACM69]. May [ACM69]. measures [Lov69]. minimal [Lov69]. minimal-program [Lov69]. models [Zei69].

non [Sav69]. non-deterministic [Sav69].

order [You69].

papers [ACM69]. partial [Ull69]. plane [Hod69]. point [Tha69]. precedence [Fis69]. Preliminary [MM69, You69]. presented [ACM69]. problem [Avi69, Ull69]. procedures [Blu69b]. processor [Kor69]. processors [Ros69]. program [Lov69]. programming [Blu69a, Zei69]. Properties [RS69, Fis69, Hod69, MP69]. Pushdown [Col69, Coo69].

Quasi [BG69]. Quasi-realtime [BG69].

real [Col69]. real-time [Col69]. realtime [BG69]. record [ACM69]. recursively [MP69]. regular [Joh69]. Report [MM69]. revisited [PR69]. Rey [ACM69].

semantics [Blu69a]. sensitive [PR69]. sets [You69]. simulation [Sav69]. Some [Fis69, Zei69]. Speed [You69]. Speed-ups [You69]. speeding [Blu69b]. stack [Ogd69]. store [Col69]. Symposium [ACM69].

theorems [Ogd69].

Theory [ACM69, ALY69, Blu69a, Tha69]. time [Avi69, Col69, Spi69]. top [RS69]. Transformations [Tha69]. Translations [AU69, Tha69]. trees [Rou69]. Turing [Sav69].

unifying [ALY69]. ups [You69].

Variations [Coo69]. Version [You69].

view [Tha69].

which [You69].

References

ACM:1969:CRA

[ACM69] ACM, editor. Conference record of ACM Symposium on Theory of Computing; papers presented at the symposium, Marina del Rey, California, May 5, 6, 7, 1969. ACM Press, New York, NY, USA, 1969. LCCN QA75.5 .A22 1969.

Amoroso:1969:UFT

[ALY69] Serafino Amoroso, Edward Lieblein, and Hisao Yamada. A unifying framework for the theory of iterative arrays of machines. In ACM [ACM69], pages 259–269. LCCN QA75.5 .A22 1969.

Aho:1969:TCF

[AU69] A. V. Aho and J. D. Ullman. Translations on a context free grammar. In ACM [ACM69], pages 93–112. LCCN QA75.5 .A22 1969.

Aviienis:1969:PCT

[Avi69] Algirdas Avižienis. On the problem of computational time and complexity of arithmetic functions. In ACM [ACM69], pages 255–258. LCCN QA75.5 .A22 1969.

Basu:1969:CCF

[Bas69] Sanat K. Basu. On classes of computable functions. In ACM [ACM69], pages 55–59. LCCN QA75.5 .A22 1969.

Book:1969:QRL

[BG69] Ronald V. Book and Sheila A.

Greibach. Quasi-realtime languages (extended abstract). In ACM [ACM69], pages 15–18. LCCN QA75.5 .A22 1969.

Blum:1969:TTS

[Blu69a] E. K. Blum. Towards a theory of semantics and compilers for programming languages. In ACM [ACM69], pages 217–227. LCCN QA75.5 .A22 1969.

Blum:1969:EPS

[Blu69b] Manuel Blum. On effective procedures for speeding up algorithms. In ACM [ACM69], pages 43–53. LCCN QA75.5 .A22 1969.

Chandler:1969:AFD

[Cha69] W. J. Chandler. Abstract families of deterministic languages. In ACM [ACM69], pages 21–30. LCCN QA75.5 .A22 1969.

Cole:1969:PSM

[Col69] Stephen N. Cole. Pushdown store machines and real-time computation. In ACM [ACM69], pages 233– 245. LCCN QA75.5 .A22 1969.

Cook:1969:VPM

[Coo69] Stephen A. Cook. Variations on pushdown machines (detailed abstract). In ACM [ACM69], pages 229–231. LCCN QA75.5 .A22 1969.

Fischer:1969:SPP

[Fis69] Michael J. Fischer. Some properties of precedence languages. In ACM [ACM69], pages 181–190. LCCN QA75.5 .A22 1969.

Herman:1969:CAF

[HI69] Gabor T. Herman and Stephen D. Isard. Computability over arbitrary fields. In ACM [ACM69], pages 149– 153. LCCN QA75.5 .A22 1969.

Hodes:1969:LCG

[Hod69] Louis Hodes. The logical complexity of geometric properties in the plane. In ACM [ACM69], pages 249–254. LCCN QA75.5 .A22 1969.

Johansen:1969:FGR

[Joh69] Peter Johansen. Free groups and regular expressions. In ACM [ACM69], pages 113–128. LCCN QA75.5 .A22 1969.

Korenjak:1969:ELP

[Kor69] A. J. Korenjak. Efficient LR(1)

processor construction. In ACM [ACM69], pages 191–200. LCCN QA75.5 .A22 1969.

Loveland:1969:MPC

[Lov69] D. W. Loveland. On minimalprogram complexity measures. In ACM [ACM69], pages 61–78. LCCN QA75.5 .A22 1969.

McCreight:1969:CCF

[MM69] E. M. McCreight and A. R. Meyer.

Classes of computable functions defined by bounds on computation: Preliminary report. In ACM [ACM69], pages 79–88. LCCN QA75.5 .A22 1969.

Manna:1969:FPR

[MP69] Zohar Manna and Amir Pnueli. Formalization of properties of recursively defined functions. In ACM [ACM69], pages 201–210. LCCN QA75.5 .A22 1969.

Ogden:1969:ITS

[Ogd69] William F. Ogden. Intercalation theorems for stack languages. In ACM [ACM69], pages 31–42. LCCN QA75.5 .A22 1969.

Peters:1969:CSI

[PR69] P. Stanley Peters and Robert W.

Ritchie. Context-sensitive immediate constituent analysis — contextfree languages revisited. In ACM [ACM69], pages 1–8. LCCN QA75.5 .A22 1969.

Rose:1969:AFP

[Ros69] Gene F. Rose. Abstract families of processors. In ACM [ACM69], pages 9–14. LCCN QA75.5 .A22 1969.

Rounds:1969:CFG

[Rou69] William C. Rounds. Contextfree grammars on trees. In ACM [ACM69], pages 143–148. LCCN QA75.5 .A22 1969.

Rosenkrantz:1969:PDT

[RS69] D. J. Rosenkrantz and R. E. Stearns. Properties of deterministic top down grammars. In ACM [ACM69], pages 165–180. LCCN QA75.5 .A22 1969.

Savitch:1969:DSN

[Sav69] Walter J. Savitch. Deterministic simulation of non-deterministic Turing machines (detailed abstract). In ACM [ACM69], pages 247–248. LCCN QA75.5 .A22 1969.

Shepard:1969:LGA

[She69] C. D. Shepard. Languages in general algebras. In ACM [ACM69], pages 155–163. LCCN QA75.5 .A22 1969.

Spira:1969:CTC

[Spi69] Philip M. Spira. On the computation time of certain classes of boolean functions. In ACM [ACM69], pages 271–272. LCCN QA75.5 .A22 1969.

Thatcher:1969:TTP

[Tha69] James W. Thatcher. Transformations and translations from the point of view of generalized finite automata theory. In ACM [ACM69], pages 129–142. LCCN QA75.5 .A22 1969.

Ullian:1969:IAP

[Ull69] Joseph S. Ullian. The inherent ambiguity partial algorithm problem for context free languages. In ACM [ACM69], pages 19–20. LCCN QA75.5 .A22 1969.

Young:1969:SUC

[You69] Paul R. Young. Speed-ups by changing the order in which sets are enumerated (preliminary version). In ACM [ACM69], pages 89–92. LCCN QA75.5 .A22 1969.

Zeiger:1969:FMS

[Zei69] H. Paul Zeiger. Formal models for some features of programming languages. In ACM [ACM69], pages 211–215. LCCN QA75.5 .A22 1969.